The following function can be expressed as the sum of two series, one of them is: x2 – 3x + 2 (A) Š(21) n=0 (B)
(C) -Σ2x) n=0 Σ(s) is0 (E) Other option
Answer #1

First the function is written in another form using partial fraction. Then using Binomial series, the power series representation is obtained. Option B is correct. All other options are incorrect. 1. is re Geven function f(n) = - 2 x²_3472 c - x(22)-1(72) x²20-142 = 234-1) Let in zaka-17 = one + = A(+1) +B1%-2) (x-2)(art